OpenStudy (anonymous):

What is the solution to the system of equations? x=4y+6 x-3y=4

OpenStudy (anonymous):

@perl

OpenStudy (mayankdevnani):

you have 2 equations :- plug 1st equation into other and solve for y ?

OpenStudy (anonymous):

is it 7 and 3 @mayankdevnani

OpenStudy (mayankdevnani):

no !!

OpenStudy (mayankdevnani):

i will tell you how to solve it

OpenStudy (mayankdevnani):

\[\large \bf x=4y+6.....(1)\] \[\large \bf x-3y=4.....(2)\]

OpenStudy (mayankdevnani):

plug 1 into 2 ,we get \[\large \bf 4y+6-3y=4\]

OpenStudy (mayankdevnani):

solve for y ??

OpenStudy (mayankdevnani):

now what is the value of y ? @igivemedalsandfollow!

OpenStudy (compphysgeek):

you have two linear equations .. draw them into the same coordinate system. you will find that the equations either are parallel and don't have a solution, they are exactly the same or they intersect in one point (x, y)